The current in a solenoid of 240 turns, having a length of 12 cm and a radius of 2 cm, changes at a rate of 0.8 A/s. Find the emf induced in it.

Grade:upto college level

1 Answers

Kevin Nash
9 years ago
Sol. E = μ base 0 N^2 A/l dl/dt = 4π x 10^-7 x (240)^2 x π(2x10^-2)^2/12 x 10^-2 x 0.8 = 4π x (24)^2 x π x 4 x 8/12 x 10^-8 = 60577.3824 x 10^-8 x 6 x 10^-4 V.